OpenStudy (anonymous):

Given the system of equations, what is the solution? x + 2y = 11 x - 2y = -1 A.){(-5, -3)} B.){(1, 1)} C.){(5, 3)}

OpenStudy (souvik):

add them and find x..

OpenStudy (uri):

x + 2y = 11 x - 2y = -1 ========= 2x=10 x=10/2 x=5

OpenStudy (anonymous):

Im not good at algebra, you tell me to find x and I have no idea how. So I don't try anymore. Sorry. But thank you ^^

OpenStudy (uri):

You want to know how did i find x?

OpenStudy (anonymous):

no, its not going to stick in my mind so I have no use to try. Im an artist, I have no use for algebra. ^^

OpenStudy (uri):

Well add both the eq so x+x=2x and +2y-2y will be cancelled and 11-1 =10.

OpenStudy (uri):

Now you have to find y.

OpenStudy (anonymous):

see....I have no idea what your talking about. im god awful at anything to do with math..

OpenStudy (uri):

x - 2y = -1 5-2y=-1(put the value of x in eq no.2) -2y=-1-5 -2y=-6 y=-6/-2 y=3